Head-First Company plans to sell 5,100 bicycle helmets at $72 each in the coming year. Variable cost is 62% of the sales price; contribution margin is 38% of the sales price. Total fixed cost equals $50,000 (includes fixed factory overhead and fixed selling and administrative expense). Required: 1. Calculate the sales revenue that Head-First must make to earn operating income of $73,120 by using the point in sales equation. 2. Check your answer by preparing a contribution margin income statement based on the sales dollars calculated in Requirement 1
Answer #1

Answer:

1. Operating income required=$73,120
Hence contribution required=(Operating income+Fixed Cost)
=(73,120+50,000)=$123,120
Selling Price per unit =$72
Contribution=(38%*72)=$27.36
Hence (27.36*units sold)=123,120
Hence units sold=4500
Sales revenue that Head-first must make=(4500*72)=$324,000
2. Contribution margin income statement
Sales= 324,000
Less:Variable cost(62%*324,000)= 200,880
Contribution= 123,120
Less:fixed cost= 50,000
Operating income =73,120